# Designing Multi-layered Runtime Guardrails for Foundation Model Based Agents: Swiss Cheese Model for AI Safety by Design

### v3

@@ Abstract
Foundation Model (FM)-based agents are revolutionizing application development across various domains. However, their rapidly growing capabilities and autonomy have raised significant concerns about AI safety. Researchers are exploring better ways to design guardrails to ensure that the runtime behavior of FM-based agents remains within specific boundaries. Nevertheless, designing effective runtime guardrails is challenging due to the agents’ autonomous and non-deterministic behavior. The involvement of multiple pipeline stages and agent artifacts, such as goals, plans, tools, at runtime further complicates these issues. Addressing these challenges at runtime requires multi-layered guardrails that operate effectively at various levels of the agent architecture. Thus, in this paper, we present a comprehensive taxonomy of runtime guardrails for FM-based agents to identify the key quality attributes for guardrails and design dimensions based on the results of a systematic literature review. Inspired by the Swiss Cheese Model, we also propose a reference architecture for designing multi-layered runtime guardrails for FM-based agents, which includes three dimensions: quality attributes, pipelines, and artifacts. The proposed taxonomy and reference architecture provide concrete and robust guidance for researchers and practitioners to build AI-safety-by-design from a software architecture perspective.
@@
Foundation Model, Large Language Models (LLM), Agent, Guardrails, Swiss Cheese Model, Responsible AI, AI Safety, Software Architecture, Taxonomy
@@ I Introduction
To address these challenges, effective runtime guardrails are key to ensure that agents behave in a safe and responsible manner [2]. In this context, guardrails are mechanisms integrated into the agent’s architecture to safeguard its behavior during runtime, preventing undesirable or unsafe behaviors [78]. There have been some initial efforts on runtime guardrails such as input filtering [9, 1], output modification [10, 11], adaptive fail-safes [12, 13], real-time monitoring and detection [14, 15, 16, 17], and continuous output validation [18, 19, 20].
@@
Therefore, in this paper, we first present a comprehensive taxonomy to categorize runtime guardrails from a software architecture perspective, based on the results of a systematic literature review. The taxonomy comprises two primary categories: quality attributes and design options. Inspired by Swiss Cheese Model [76], we also propose novel reference architecture for designing multi-layered guardrails of FM-based agents which include three dimensions: quality attributes, pipelines, and artifacts. Each guardrail layer can be designed to protect specific quality attributes (such as privacy and security), specific pipeline stages (such as prompts, intermediate results and final results), as well as agent artifacts (such as goals, plans, and tools). While each layer may have its own weaknesses (i.e. holes in the Swiss Cheese Model), the combined layers create a a robust defense against failures. This reference architecture provides concrete guidance for researchers and practitioners, enabling AI-safety-by-design from a software architecture perspective.
The rest of the paper is organized as follows. section II discuss the related works and background study required to understand the proposed works. The research methods employed in this study are described in section III, including a brief discussion of the research protocol used for systematic literature review. The proposed taxonomy of guardrails is presented in section IV,developed based on the results of a systematic literature review. The taxonomy is organized into guardrails quality attributes and design options from different perspectives. section V proposes the reference architecture for multi-layered runtime guardrails for FM-based agents. section VI identifies and summarizes the primary threats that could impact the validity of this study. Finally, section VII concludes the paper and outlines directions for future work.
Figure 1: Methodology
@@ II Background and Related Work
FMs have significantly advanced current agent development and emphasize the need to safeguard their behavior [3, 20]. In this context, guardrails for FM-based agents have been explored; however, there is a lack of comprehensive studies that provide a thorough understanding of guardrails for FM-based agents. This paper aims to fill this gap. In the following sections, we present key background information and related work.
@@ II-A Recent State-of-the-Art Works on Foundation Models and FM-Based Agents
In 2021, Bommasani et al. [1] provided a comprehensive discussion on FMs, illustrating key elements, relationships, opportunities, and associated risks. While their focus was on FMs in general, they highlighted the potential for these models to serve as the foundation for more complex systems, including FM-based agents. Zhou et al. [20] reviewed research advancements, challenges, and opportunities for pre-trained models in text, image, graph, and data modalities. They also discussed the integration of FMs into systems such as agents. Both works offer excellent insights into future research directions to address open problems and associated risks in FM-based agents.
Recently, Lu et al. developed a taxonomy of FM-based systems focusing on their pre-training, adaptation, architectural design, and responsible-AI-by-design [27]. The taxonomy aids software architects and developers in evaluating and integrating FMs into complex agent systems. The authors then highlighted considerations for responsible AI and safety attributes. Several other works [30, 2, 31, 36, 8], also emphasize the importance of responsible AI and safety practices for FM-based agents. In [28, 29], the authors explored the risks associated with deploying LLM-based agents and evaluated current approaches for mitigating these risks through model alignment, respectively. In 2024, a reference architecture for designing responsible and safe FM-based agents is proposed in [3]. The authors demonstrated that the unique characteristics of FM-based agents—such as their autonomous operation, non-deterministic behavior, and continuous evolution—pose significant challenges in ensuring responsible AI and AI safety.
II-B Existing Guardrails Approaches and Tools for FM-Based Agents
@@ III Methodology
This study focuses on two primary concepts: (i) foundation model-based agents and (ii) runtime guardrails. We adopted the Petticrew and Roberts approach [39] to define the Population, Interventions, Comparison, Outcomes, and Context (PICOC), within which the intervention in this study is delivered. The PICOC for this study is shown in Table I. Using these PICOC components and following Kitchenham’s guidelines [40], we develop the protocol for this study.
@@
III-A Research Scope and Protocol Development
The high-level research approach for this study is shown in Figure 1. Initially, we determined the research scope and developed a protocol following Kitchenham’s guidelines [40, 41]. The protocol guided the entire study by defining relevant scientific databases and resources, formulating keywords and search strings, outlining qualitative and quantitative checklists, and specify criteria for study inclusion and exclusion.
III-B Research Questions
Our first research question studies the key qualities for designing multi-layered runtime guardrails in FM-based agents. section IV-A elaborate on how this research question is addressed.
Our second research question investigates guardrails design options in FM-based agents from different perspectives, including action, target, scope, rule, autonomy, modalities, and underlying techniques. section IV-B outlines our approach to addressing this research question.
Our third research question explores how to address the challenges arising from the autonomous and deterministic nature of FM-based agents. Specially, we examine how to adapt the Swiss Cheese Model to safeguard the behaviors of FM-based agents by implementing multi-layered guardrails across various agent artifacts. section V presents the proposed architecture and discusses our strategies for addressing this research question.
TABLE II: Consolidated Concepts and Search Terms
Main Terms Supportive Search Terms
Concept 1 (Co1): Foundation Model based agents Foundation Models, Foundation Model based agents, Large Language Model, Generative AI, Artificial General Intelligence, Transformer Models, Self-supervised Learning, Pretrained Models, Language Models, Conversational AI.
Concept 2 (Co2): Runtime Guardrails Guardrails, guardian, responsible AI, safe, risk, trustworthy, protect, detect, monitor, verify, validate, evaluate, benchmark, design.
III-C Search String Formulation
Relevant primary studies for this SLR were identified based on the RQs defined in section III-B. With the assistance of the PICOC approach (shown in Table I), our search terms were divided into two primary concepts, as shown in Table II. These concepts helped us to set a well-formulated search string.
[…]
[$[\{(C_{11}\|C_{12}\|...\|C_{1n})\textbf{AND}(C_{21}\|C_{22}\|...\|C_{2n})$ (1)
$\textbf{NOT}(UC_{1}\|UC_{2}\|...\|UC_{n})]$
where $C_{11,12,...,1n}$, and $C_{21,22,...,2n}$ $\varepsilon$ Co1 and Co2 of Table II, respectively; and $UC_{1}$, $UC_{2}$, . . ., $UC_{n}$ refers the Exclude Context defined earlier in PICOC (Table I).
III-D Selection of Papers: Inclusion and Exclusion Criterion
Table III and Table IV present the Inclusion Criteria (IC) and Exclusion Criteria (EC) that have been used to identify the studies for this SLR, respectively. We found that a considerable amount of work on guardrails exists in gray literature; however, we excluded them as they often lack peer review and a rigorous validation process. While some sources [42] argue that gray literature is an important resource for systematic literature reviews (SLRs), such literature can be misleading and introduce biases and inconsistencies in the review process [43]. We prioritized peer-reviewed sources in this study to ensure scientific reliability and credibility, as per Kitchenham et al. guidelines [40, 41].
TABLE III: Inclusion Criteria
ID Detail Criterion
IC1 Full text of conference papers, journal articles, industry reports, and book chapters that are relevant to the defined main concepts: Foundation model based agents and guardrails.
IC2 Papers written in English that include references.
IC3 Studies that specifically address the design and development of guardrails in foundation model-based agents. This includes theoretical frameworks, empirical research and case studies.
IC4 Papers available in an electronic format, such as PDF, DOC, DOCX, HTML, and PS etc.
TABLE IV: Exclusion Criteria
ID Detail Criterion
EC1 Work-in-progress proposals, keynote addresses, secondary studies, and vision papers without concrete relation to guardrails.
EC2 Discussion papers and opinion pieces that do not provide empirical evidence or concrete solutions related to guardrails in foundation model-based agents.
EC3 Short communications less than two pages, and studies that do not offer substantial information for analysis.
EC4 Studies focusing solely on AI or similar technologies without direct relevance to guardrails.
EC5 Research lacking a clear connection to the design and development of guardrails in the context of foundation models.
EC6 Duplicate publications or earlier versions of studies that have been superseded by extended journal versions.
EC7 Non-original research, commentary, editorial pieces, and non-empirical discussions papers.
EC8 Studies inaccessible due to copyright or database restrictions.
III-E Study Search and Filtering Process
Our filtration process is further detailed in Figure 2. Initially we ran the formatted query on four major databases that returned 1,733 research papers. We then applied filtering and classified the studies found according to the guidelines presented in [41, 40]. In our initial filtration process, we removed 108 papers due to being duplicated articles, editorial or key notes. After reading the title, abstract, conclusion and skimming through the introduction, methodology and results, we applied our exclusion criterion defined in Table IV, and 1524 further papers were removed. During the third step of filtration, we applied inclusion criteria and removed 80 papers as these studies did not meet ICs shown in Table IV. In parallel, we did a manual search and found 189 papers that meet our key concepts defined in Table II but not contain any unwanted content (UC). After applying ICs and ECs, 15 out of 189 papers were selected. Finally, we did a cross-check and ended up with 32 papers (shown in Appendix A).
Figure 2: Study Selection Process for this SLR
III-F Data Extraction and Quality Assessment
@@
We used a semi-automated process [44] for data extraction from the selected studies to answer our RQs. Key qualitative information extracted from each selected study includes guardrails definitions, motivations, reported key quality attributes, and design options. We also extracted several relevant pieces of information to understand the context and considerations in designing and evaluating runtime guardrails.
@@
We then evaluated each study based on the following five Quality Assessment Criteria (QAC) on a scale from 1 (Very Poor) to 5 (Excellent). If a study’s average score was less than 2, it was excluded from further analysis. Otherwise, we used the qualitative information to decide this. The QAC used for this study are  [^1] QAC score for each selected study is presented in Appendix B:
  • Relevance to guardrails for FM-based agents.
  • Clear methodology for guardrail design.
  • Adequate data collection, analysis, and evaluation of guardrail effectiveness at different layers of the agent architecture.
  • Discussion of challenges in designing guardrails for autonomous and non-deterministic behaviors in agents.
  • Practical applicability of findings for guardrails in FM-based agents.
@@
IV Taxonomy of Guardrails for FM-based Agents
Figure 3 presents the proposed taxonomy of runtime guardrails for FM-based agents, developed based on the results of a systematic literature review. The taxonomy is organized into external and internal quality attributes, and design options from different perspectives.
IV-A Quality Attributes of Guardrails
We examine the key quality attributes that should be considered when designing runtime guardrails. Below, we discuss these attributes in detail.
IV-A1 Accuracy
Accuracy in FM-based agents is crucial, particularly in mitigating issues such as hallucinations, misinformation, and disinformation [45]. Hallucinations occur when models generate information that is factually incorrect. Such inaccuracies can mislead users and damage the credibility of the agent [10]. Misinformation refers to the unintentional spread of false information, while disinformation involves the deliberate dissemination of falsehoods to receive users [20]. For example, OpenAI uses guardrails to clearly label AI-generated content to prevent deepfakes and misinformation [46]. One such case has been reported to prevent misleading voters in last US elections [47, 48].
IV-A2 Efficiency
Efficiency is crucial in FM-based agents, as users expect fast, efficient responses [24]. Without guardrails, agents risk engaging in resource-intensive tasks that slow down response times [54]. By dynamically managing resources across multiple layers, these guardrails prevent inefficiencies, such as endless loops, and filter irrelevant inputs, ensuring that agents focus on processing meaningful data [64, 16, 3]. Additionally, FM-based agents can incur significant costs due to errors, inefficiencies, or non-compliance with regulations. Without proper guardrails, agents might generate outputs that lead to financial losses, legal penalties, or damage to their reputation [26, 12]. For example, an agent that provides incorrect financial advice could result in monetary losses for users and potential lawsuits against the provider.
@@ IV-A3 Privacy
Privacy in FM-based agents poses risks due to handling sensitive data, where data leakage might expose personal information [12, 1]. This leakage can occur through direct responses or statistical inferences, or inadvertent revelations through model outputs. In April-May 2023, a notable incident involved Samsung employees leaking proprietary information into ChatGPT, leading to Samsung banning ChatGPT [50].
@@ IV-A7 Compliance
Compliance in FM-based agents involves adhering to legal and regulatory standards [16, 20]. These issues are critical because non-compliance can lead to legal penalties, reputational damage, and loss of user trust. Runtime guardrails reduce these risks by ensuring alignment with data protection regulations, industry standards, and guidelines through continuous monitoring at multiple levels [54, 26]. Additionally, these guardrails assist in automating compliance checks. They ensure that all aspects of the FM-based agent’s operations align with the necessary legal and regulatory frameworks [62, 36], and better support internal audits and external regulatory reviews [12]. For example, FM-based agents may unintentionally facilitate unauthorized use of generated content, making it vulnerable to duplication or improper distribution [26, 55, 10]. Guardrails operating in real time help mitigate these risks by detecting and restricting unauthorized access, ensuring better copyright protection [49]. Techniques such as watermarking, fingerprinting, and labeling are applied across different layers to ensure the ownership and compliance with licensing laws [10, 1].

### v4

# Swiss Cheese Model for AI Safety: A Taxonomy and Reference Architecture for Multi-Layered Guardrails of Foundation Model Based Agents

@@ Abstract
Foundation Model (FM)-based agents are revolutionizing application development across various domains. However, their rapidly growing capabilities and autonomy have raised significant concerns about AI safety. Researchers are exploring better ways to design guardrails to ensure that the runtime behavior of FM-based agents remains within specific boundaries. Nevertheless, designing effective runtime guardrails is challenging due to the agents’ autonomous and non-deterministic behavior. The involvement of multiple pipeline stages and agent artifacts, such as goals, plans, tools, at runtime further complicates these issues. Addressing these challenges at runtime requires multi-layered guardrails that operate effectively at various levels of the agent architecture. Therefore, in this paper, based on the results of a systematic literature review, we present a comprehensive taxonomy of runtime guardrails for FM-based agents to identify the key quality attributes for guardrails and design dimensions. Inspired by the Swiss Cheese Model, we also propose a reference architecture for designing multi-layered runtime guardrails for FM-based agents, which includes three dimensions: quality attributes, pipelines, and artifacts. The proposed taxonomy and reference architecture provide concrete and robust guidance for researchers and practitioners to build AI-safety-by-design from a software architecture perspective.
@@
Foundation Model, Large Language Models, LLM, Agent, Guardrails, Safeguard, AI Safety, Software Architecture, Taxonomy, Swiss Cheese Model, Responsible AI
[…]
@@ I Introduction
To address these challenges, effective runtime guardrails are key to ensure that agents behave in a safe and responsible manner [3]. In this context, guardrails are mechanisms integrated into an agent’s architecture to safeguard its behavior during runtime, preventing undesirable or unsafe behaviors [4]. There have been some initial efforts on runtime guardrails such as input filtering [6, 1], output modification [7, 8], adaptive fail-safes [9, 10], real-time monitoring and detection [11, 12, 13, 14], and continuous output validation [15, 16, 17].
@@
Therefore, in this paper, we first present a comprehensive taxonomy to categorize runtime guardrails from a software architecture perspective, based on the results of a systematic literature review. The taxonomy comprises two primary categories: quality attributes and design options. Inspired by Swiss Cheese Model [18], we also propose novel reference architecture for designing multi-layered guardrails of FM-based agents which include three dimensions: quality attributes, pipelines, and artifacts. Each guardrail layer can be designed to protect specific quality attributes (such as privacy and security), specific pipeline stages (such as prompts, intermediate results and final results), as well as agent artifacts (such as goals, plans, and tools). While each layer may have its own weaknesses (i.e. holes in the Swiss Cheese Model), the combined layers create a robust defense against failures. This reference architecture provides concrete guidance for researchers and practitioners, enabling AI-safety-by-design from a software architecture perspective.
The rest of the paper is organized as follows. section II discuss the related works. The research methods employed in this study are described in section III. The proposed taxonomy of guardrails is presented in section IV. section V proposes the reference architecture for multi-layered runtime guardrails for FM-based agents. section VI discusses threats, and section VII concludes the paper with future work.
@@ II Background and Related Work
FMs have significantly advanced current agent development and emphasize the need to safeguard their behavior [2, 17]. In this context, guardrails for FM-based agents have been explored; however, but comprehensive studies on this topic are lacking. This paper addresses that gap.
@@ II-A Recent State-of-the-Art Works on Foundation Models and FM-Based Agents
In 2021, Bommasani et al. [1] provided a comprehensive discussion on FMs, illustrating key elements, relationships, opportunities, and associated risks. While their focus was on FMs in general, they highlighted the potential for these models to serve as the foundation for more complex systems, including FM-based agents. Zhou et al. [17] reviewed research advancements, challenges, and opportunities for pre-trained models in text, image, graph, and data modalities. They also discussed the integration of FMs into systems such as agents. Both works offer excellent insights into future research directions to address open problems and associated risks.
Recently, Lu et al. developed a taxonomy of FM-based systems focusing on their pre-training, adaptation, architectural design, and responsible-AI-by-design [19]. The taxonomy aids software architects and developers in evaluating and integrating FMs into complex agent systems. The authors then highlighted considerations for responsible AI and safety attributes. Several other works [20, 3, 21, 22, 23] also emphasize the importance of responsible AI and safety practices for FM-based agents. In [24, 25], the authors explored the risks associated with deploying LLM-based agents and evaluated current approaches for mitigating these risks through model alignment, respectively. In 2024, a reference architecture for designing responsible and safe FM-based agents is proposed in [2]. The authors demonstrated that the unique characteristics of FM-based agents—such as their autonomous operation, non-deterministic behavior, and continuous evolution—pose significant challenges in ensuring responsible AI and AI safety.
II-B Guardrails Approaches and Tools for FM-Based Agents
@@ III Methodology
This study focuses on two primary concepts: (i) FM-based agents and (ii) multi-layered runtime guardrails. We adopted the Petticrew and Roberts approach [33] to define the Population, Interventions, Comparison, Outcomes, and Context (PICOC), within which the intervention in this study is delivered. The PICOC for this study is shown in Table I. Using these PICOC components and following Kitchenham’s guidelines [34], we developed the protocol for this study.
III-A Research Approach
The research approach for this study is summarized in Figure 1. Initially, we determined the research scope and developed a protocol following Kitchenham’s guidelines [34]. The protocol guided the entire study by defining relevant scientific resources, formulating keywords and search strings, outlining qualitative and quantitative checklists, and specifying criteria for the inclusion and exclusion of studies [^1] A detailed review protocol discussion is beyond this paper’s scope; see supplementary materials for details [35]..
Figure 1: Methodology
III-B Research Questions
Our first research question studies the key qualities for designing multi-layered runtime guardrails in FM-based agents. subsection IV-A elaborate on how it is addressed.
Our second research question investigates guardrails design options in FM-based agents from different perspectives, including action, target, scope, rule, autonomy, modalities, and underlying techniques. subsection IV-B outlines our approach to addressing this research question.
Our third research question explores how to address the challenges arising from the autonomous and non-deterministic nature of FM-based agents. Specially, we examine how to adapt the Swiss Cheese Model to safeguard the behaviors of FM-based agents by implementing multi-layered guardrails across various agent artifacts. section V presents the proposed architecture and discusses our strategies for addressing this research question.
@@ III-B Research Questions
III-C Study Search and Filtering Process
To collect potentially relevant studies, we executed an automated keyword-based search and carried out the study filtering process in March 2024. Consequently, studies published after this period were not investigated in this research. Our filtration process involved initial screening, grouping papers, and resolving conflicts in study selection. We also conducted manual searching and backward snowballing to ensure no significant studies were missed. The study filtering process, including the list of selected studies and their quality assessment scores, is detailed in supplementary materials [35]. We have also made our study filtration file (‘Study Filtration.xlsx’) publicly available in the supplementary materials. This file presents a step-by-step breakdown of the filtration process, narrowing the initially retrieved studies to the final selected studies. Once all relevant studies were cross-checked, identified, and collected, we performed data extraction, synthesis, and analysis. Finally, findings from 32 high-quality selected studies were used to develop the taxonomy and presented in this report.
III-D Data Extraction and Quality Assessment
@@
We used a semi-automated process [36] for data extraction from the selected studies to answer our RQs. Key qualitative information extracted from each selected study includes guardrails definitions, key quality attributes, and design options. We also extracted several relevant pieces of information to understand the context and considerations in designing and evaluating multi-layered runtime guardrails [^2] Data extraction sheet is provided in the supplementary materials [35].
@@
We evaluated each study based on following five Quality Assessment Criteria (QAC) on a scale from 1 (Very Poor) to 5 (Excellent): (i) relevance to guardrails for FM-based agents, (ii) clarity of methodology for guardrail design, (iii) adequacy in data collection, analysis, and evaluation of guardrail effectiveness across different layers of the agent architecture, (iv) discussion of challenges in designing guardrails for autonomous and non-deterministic behaviors in agents, and (v) practical applicability of findings for guardrails in FM-based agents. If a study’s average score was less than 2, it was excluded from further analysis. Otherwise, we used the qualitative information to decide this.
IV Taxonomy of Guardrails for FM-based Agents
Figure 2 presents the proposed taxonomy of runtime guardrails for FM-based agents, developed based on the results of a systematic literature review. The taxonomy was developed by synthesizing findings from 32 selected studies through an iterative clustering and validation process. It is organized into two key components: (i) Quality attributes, capturing relevant facets that influence guardrails’ effectiveness and reliability, and (ii) Design dimensions, representing practical approaches for implementing guardrails. The taxonomy combines external and internal quality attributes to merge overlapping categories, as some attributes (e.g., security, adaptability) apply to both developer-focused and stakeholder-facing contexts. Here, we also standardize terminology and refine overlapping attributes based on established practices in software engineering [34, 37], while acknowledging interdependencies. This organization helps researchers and developers to distinguish between guardrails’ core characteristics from practical implementation approaches.
IV-A Quality Attributes of Guardrails
We examine the key quality attributes essential for designing runtime guardrails. These attributes ensure guardrails meet critical performance, security, and reliability goals aligned with design objectives, while also improving end-user trust and system-level outcomes. They were systematically synthesized, clustered, and refined through an iterative process to emphasize their unique contributions. Below, we discuss these attributes in detail.
IV-A1 Accuracy
Accuracy in FM-based agents is crucial, particularly in mitigating issues such as hallucinations, misinformation, and disinformation [38]. Hallucinations occur when models generate information that is factually incorrect. Such inaccuracies can mislead users and damage the credibility of the agent [7]. Misinformation refers to the unintentional spread of false information, while disinformation involves the deliberate dissemination of falsehoods to deceive users [17]. FM-based agents can incur significant costs due to errors, inaccuracies (leading to inefficiency), or non-compliance with regulations. Without proper guardrails, agents might generate outputs that lead to financial losses, legal penalties, or damage to their reputation [39, 9]. Thus, leading companies e.g., OpenAI and Google use guardrails to clearly label AI-generated content to prevent deepfakes and misinformation [40, 41].
IV-A2 Efficiency
Efficiency is crucial in FM-based agents, as users expect fast, efficient responses [42]. Without guardrails, agents risk engaging in resource-intensive tasks that slow down response times [43]. By dynamically managing resources across multiple layers, these guardrails prevent inefficiencies, such as endless loops, and filter irrelevant inputs, ensuring that agents focus on processing meaningful data [44, 13, 2].
@@ IV-A3 Privacy
Privacy in FM-based agents poses risks due to handling sensitive data, where data leakage might expose confidential information [9, 1]. This leakage can occur through direct responses or statistical inferences, or inadvertent revelations through model outputs. In 2023, a notable incident involved Samsung employees leaking proprietary information into ChatGPT, leading to Samsung banning ChatGPT [45].
@@ IV-A7 Compliance
Compliance in FM-based agents involves adhering to legal and regulatory standards [13, 17]. These issues are critical because non-compliance can lead to legal penalties, reputational damage, and loss of user trust. Runtime guardrails reduce these risks by ensuring alignment with data protection regulations, industry standards, and guidelines through continuous monitoring at multiple levels [43, 39]. Additionally, these guardrails assist in automating compliance checks. They ensure that all aspects of the FM-based agent’s operations align with the necessary legal and regulatory frameworks [29, 22], and better support internal audits and external reviews [9]. For example, FM-based agents may unintentionally facilitate unauthorized use of generated content, making it vulnerable to duplication or improper distribution [39, 26, 7]. Guardrails operating in real time help mitigate these risks by detecting and restricting unauthorized access, ensuring better copyright protection [50]. Techniques such as watermarking and labeling are applied across different layers to ensure the ownership and compliance with licensing laws [7, 1].